# Classical and Quantum KMS States on Spin Lattice Systems

### v2

@@ 1 Introduction
Strict (or $C^{*}$-algebraic) deformation quantization (SDQ) provides a mathematically rigorous setting to study the quantization of a classical system [36, 48]. This framework is not only suitable to investigate the semiclassical limit of states of a quantum system with a fixed, but arbitrary number of degrees of freedom cf. [37, 40], but it can also be applied to describe the macroscopic properties of quantum systems over an infinitely extended lattice [20, 38, 39, 52, 53].
From a mathematical point of view, a SDQ requires the notion of bundle of $C^{*}$-algebra which we briefly recall following [37, App. C.19], [10, §IV.1.6]. Setting $\overline{\mathbb{Z}_{+}/2}:=\mathbb{Z}_{+}/2\cup\{\infty\}$ and given a collection $\{\mathfrak{A}_{j}\}_{j\in\overline{\mathbb{Z}_{+}/2}}$ of $C^{*}$-algebras, we denote by $\prod_{j\in\overline{\mathbb{Z}_{+}/2}}\mathfrak{A}_{j}$ the associated full $C^{*}$-direct product, which is the $C^{*}$-algebra made by sequences $(\mathfrak{a}_{j})_{j\in\overline{\mathbb{Z}_{+}/2}}$, $\mathfrak{a}_{j}\in\mathfrak{A}_{j}$, such that $\|(\mathfrak{a}_{j})_{j\in\overline{\mathbb{Z}_{+}/2}}\|_{\prod_{j\in\overline{\mathbb{Z}_{+}/2}}\mathfrak{A}_{j}}:=\sup_{j}\|\mathfrak{a}_{j}\|_{\mathfrak{A}_{j}}<\infty$. Within this setting a continuous bundle of $C^{*}$-algebras over $\overline{\mathbb{Z}_{+}/2}$ (with fibers $\{\mathfrak{A}_{j}\}_{j\in\overline{\mathbb{Z}_{+}/2}}$) is a $C^{*}$-subalgebra $\mathfrak{A}\subset\prod_{j\in\overline{\mathbb{Z}_{+}/2}}\mathfrak{A}_{j}$ such that: (i) $(\|\mathfrak{a}_{j}\|_{j\in\overline{\mathbb{Z}_{+}/2}})\in C(\overline{\mathbb{Z}_{+}/2})$ for all $(\mathfrak{a}_{j})_{j\in\overline{\mathbb{Z}_{+}/2}}\in\mathfrak{A}$, where $C(\overline{\mathbb{Z}_{+}/2})$ denotes the space of sequences $(\alpha_{j})_{j\in\overline{\mathbb{Z}_{+}/2}}$, $\alpha_{j}\in\mathbb{C}$, such that $\alpha_{\infty}=\lim_{j\to\infty}\alpha_{j}$; (ii) $\alpha\mathfrak{a}\in\mathfrak{A}$ for all $\mathfrak{a}\in\mathfrak{A}$ and $\alpha\in C(\overline{\mathbb{Z}_{+}/2})$. A strict deformation quantization (SDQ) is then defined by the following data:
@@
    We investigate further the relationship between classical and quantum thermal equilibrium with a specific focus on phase transitions. The latter describe the uniqueness/non-uniqueness of KMS states and are of utmost relevance for describing when a physical system undergoes an abrupt change in its macroscopic behaviour, e.g. gas-to-liquid condensation. It is common folklore that classical and quantum phase transition at non-vanishing temperature should be in bijection. Yet, to the best of our knowledge, no mathematically rigorous proof of this claim has been given. In Section 4 we prove that, within the model $\mathbb{S}^{2}$-valued lattice spin system considered in Section 2, under a sufficiently mild assumption, cf. (44), classical and quantum KMS states are unique for temperatures higher than a common sufficiently high threshold temperature. This results is a consequence of Theorems 4.1, 4.7 which provides two new sufficient conditions for the uniqueness of KMS classical and quantum states. Our result provides a first rigorous indication of the equivalence between classical and quantum phase transitions.
It is worth to point out that our results are companions of other existing works in this area. In particular, [41] already described an abstract framework which covers the Berezin SDQ for a lattice system in a finite region. Our result (I) generalizes this setting for a specific model but allowing to deal with a spin lattice system on an infinitely extended region: This is important because physically interesting results on thermal equilibrium, e.g. phase transitions, can only be described on infinitely extended system. Concerning (II), in [23] the semiclassical limit of KMS quantum states has been investigated under an assumption which is an abstract version of our Lemma 3.4. Similarly, [54] deals with the semiclassical limit of Gibbs quantum and classical states, i.e. KMS states on a spin lattice system associated with a finite region: From this point of view, our result (II) can be seen as a generalization of [54] to a physically more interesting scenario. Finally, concerning (III), Theorems 4.1, 4.7 are inspired by [13, Prop. 6.2.45], which provides a sufficient condition for uniqueness of KMS quantum states with an argument based on a quantum version of the Kirkwood-Salzburg equations. Our result provides a classical analogous of [13, Prop. 6.2.45], moreover, it strengths some of its conclusion —specifically the $j$-dependence of the inverse critical temperature, cf. Remark 4.9.
@@ 2.2 Berezin SDQ for a single site system
To begin with we consider the Lie group $SU(2)$ and denote by $\{J_{i}\}_{i=1}^{3}$ the generators of the corresponding Lie algebra $\mathfrak{su}(2)$ with commutation relations $[J_{1},J_{2}]=iJ_{3}$ and extended cyclically. Adopting the standard physicist’s notation we denote by $\boxed{D^{(j)}}\colon SU(2)\to M_{2j+1}(\mathbb{C})$ the irreducible representation of $SU(2)$ of spin $j\in\mathbb{Z}_{+}/2$, cf. [27, §5.4]. We will denote by
$|j,m\rangle\in\mathbb{C}^{2j+1}\qquad m\in[-j,j]\cap\mathbb{Z}\,,$ (10)
@@
$Q_{j}(a):=\int_{\mathbb{S}^{2}}a(\sigma)|j,m\rangle\langle j,m|\,\mathrm{d}\mu_{j}(\sigma)\,.$ (13)
@@
identify a SDQ over a suitably defined bundle of $C^{*}$-algebras $\boxed{B_{\ast}}\subseteq\prod_{j\in\overline{\mathbb{Z}_{+}/2}}B_{j}$, cf. [37, Thm. 8.1]. From a physical point of view, the semiclassical parameter is identified with $\boxed{h_{j}}:=(2j+1)^{-1}$.
@@
where $P_{\ell,m}$ denotes the Legendre polynomial of order $\ell,m$. Notice that this choice of normalization is such that $\|Y_{\ell,m}\|_{L^{2}(\mathbb{S}^{2},\mu_{\ell})}=1$, moreover, $\|Y_{\ell,m}\|_{B_{\infty}}\leq 1$, cf. [51, Cor. 2.9]. The set $\{Y_{\ell,m}\}_{\ell,m}$ is a complete orthogonal system for $L^{2}(\mathbb{S}^{2},\mu_{0})$ made by orthogonal but not $L^{2}$-normalized vectors. With this convention we also have [42, §3.6.2]
@@
where we used Schur orthogonality relations, cf. [34, §4.10], while $\boxed{\textsc{cg}_{j_{1},m_{1};j_{2},m_{2}}^{j,m}}$ denotes the Clebsch-Gordan coefficients, cf. [42, §3]. We recall in particular that $\textsc{cg}_{j_{1},m_{1};j_{2},m_{2}}^{j,m}\in\mathbb{R}$, moreover, the coefficient vanishes unless $m=m_{1}+m_{2}$ and $|j_{1}-j_{2}|\leq j\leq j_{1}+j_{2}$. For later convenience it is also worth recalling that, for all $\sigma\in\mathbb{S}^{2}$,
@@
$\langle A|B\rangle_{\textsc{hs}}:=\frac{1}{j+1}\operatorname{tr}(A^{*}B)\,,$ (19)
@@
By direct inspection $\tilde{D}^{(j)}(R)$ is unitary with respect to the Hilbert-Schmidt scalar product (19) on $M_{2j+1}(\mathbb{C})$, moreover $R\mapsto\tilde{D}^{(j)}(R)$ is a unitary representation of $SU(2)$. Though $D^{(j)}$ is irreducible, $\tilde{D}^{(j)}$ is not irreducible and by Peter-Weyl Theorem, cf. [27, §5.2], it decomposes into irredubile representations $\{D^{(\ell)}\}_{\ell\in\mathbb{Z}_{+}/2}$. Notably the Berezin quantization map $Q_{j}\colon C(\mathbb{S}^{2})\to M_{2j+1}(\mathbb{C})$ of Equation (13) can be used to provide the explicit decomposition of $\tilde{D}^{(j)}$ in its irreducible components. To this avail, let
@@
where in the last line we used the symmetry property of the Clebsch-Gordan coefficients, cf. [34, §8]. Equations (23)-(25) imply that
@@
form a complete orthogonal system in $M_{2j+1}(\mathbb{C})$ with respect to the Hilbert-Schmidt scalar product (19). (As an aside, we observe that this fact provides a quick proof of that $Q_{j}\colon\dot{B}_{\infty}\to B_{j}$ is surjective: Indeed, for any $A_{j}\in B_{j}$ we may consider $a_{j}:=\sum_{\ell=0}^{2j}\sum_{m=-\ell}^{\ell}A_{\ell,m}Y_{\ell,m}\in\dot{B}_{\infty}$ where $A_{\ell,m}:=\|Q_{j}(Y_{\ell,m})\|_{\textsc{hs}}^{-2}\langle Q_{j}(Y_{\ell,m})|A_{j}\rangle_{\textsc{hs}}$ so that $A_{j}=Q_{j}(a_{j})$.) Moreover, Equation (22) ensures that $Q_{j}$ intertwines between the left-action representation $R\mapsto\hat{R}$ and $\tilde{D}^{(j)}$. Since $R\mapsto\hat{R}$ is unitary equivalent to $D^{(\ell)}$ when restricted to the vector space spanned by $\{Y_{\ell,m}\}_{m\in[-\ell,\ell]\cap\mathbb{Z}}$, it follows that $\tilde{D}^{(j)}$ is unitary equivalent to $D^{(\ell)}$ when restricted to the vector space spanned by $\{Q_{j}(Y_{\ell,m})\}_{m\in[-\ell,\ell]\cap\mathbb{Z}}$. Thus, $\tilde{D}^{(j)}$ is not irreducible and decomposes into direct sum of the irreducibile representations of $SU(2)$ with total spin $\ell\in\{0,\ldots,2j\}$, each of which taken with multiplicity one.
@@ 2.3 Berezin SDQ on $\Gamma$
To this avail we recall that in [41] the Berezin SDQ $Q_{j}\colon\dot{B}_{\infty}\to B_{j}$ has been lifted to a SDQ $\boxed{Q_{j}^{\Lambda}}\colon\dot{B}_{\infty}^{\Lambda}\to B_{j}^{\Lambda}$ for any finite region $\Lambda\Subset\Gamma$. In a nutshell, this boils down to define $Q_{j}^{\Lambda}$ by linear extension of the tensor product map $\bigotimes_{x\in\Lambda}Q_{j}$ and checking that the data $B_{\infty}^{\Gamma},B_{\ast}^{\Lambda},\{Q_{j}\}_{j\in\mathbb{Z}_{+}/2}$ fulfil the requirement of a SDQ —here $\boxed{B_{\ast}^{\Lambda}}:=\bigotimes_{x\in\Lambda}B_{\ast}$. A non-trivial task in this setting is to prove that $B_{\ast}^{\Lambda}$ is again a bundle of $C^{*}$-algebras over $\prod_{j\in\overline{\mathbb{Z}_{+}/2}}B_{j}^{\Lambda}$: This is addressed in full generality in [41] by using the results of [33].
@@
Let $\{\mathfrak{A}_{j}\}_{j\in\overline{\mathbb{Z}_{+}/2}}$ be a collection of $C^{*}$-algebras. For later convenience we recall the following sufficient condition which identifies a continuous bundle of $C^{*}$-algebras $\mathfrak{A}\subset\prod_{j\in\overline{\mathbb{Z}_{+}/2}}\mathfrak{A}_{j}$ by defining a dense set of (a posteriori) elements of $\mathfrak{A}$cf. [36, Prop. 1.2.3], [37, Prop. C.124]. In more details, let $\dot{\mathfrak{A}}\subseteq\prod_{j\in\overline{\mathbb{Z}_{+}/2}}\mathfrak{A}_{j}$ be such that:

### v3

@@ 1 Introduction
Strict (or $C^{*}$-algebraic) deformation quantization (SDQ) provides a mathematically rigorous setting to study the quantization of a classical system [39, 52]. This framework is not only suitable to investigate the semiclassical limit of states of a quantum system with a fixed, but arbitrary number of degrees of freedom cf. [40, 43], but it can also be applied to describe the macroscopic properties of quantum systems over an infinitely extended lattice [22, 41, 42, 56, 57] or the semi-classical properties of condensates [48].
From a mathematical point of view, a SDQ requires the notion of bundle of $C^{*}$-algebra which we briefly recall following [40, App. C.19], [12, §IV.1.6]. Setting $\overline{\mathbb{Z}_{+}/2}:=\mathbb{Z}_{+}/2\cup\{\infty\}$ and given a collection $\{\mathfrak{A}_{j}\}_{j\in\overline{\mathbb{Z}_{+}/2}}$ of $C^{*}$-algebras, we denote by $\prod_{j\in\overline{\mathbb{Z}_{+}/2}}\mathfrak{A}_{j}$ the associated full $C^{*}$-direct product, which is the $C^{*}$-algebra made by sequences $(\mathfrak{a}_{j})_{j\in\overline{\mathbb{Z}_{+}/2}}$, $\mathfrak{a}_{j}\in\mathfrak{A}_{j}$, such that $\|(\mathfrak{a}_{j})_{j\in\overline{\mathbb{Z}_{+}/2}}\|_{\prod_{j\in\overline{\mathbb{Z}_{+}/2}}\mathfrak{A}_{j}}:=\sup_{j}\|\mathfrak{a}_{j}\|_{\mathfrak{A}_{j}}<\infty$. Within this setting a continuous bundle of $C^{*}$-algebras over $\overline{\mathbb{Z}_{+}/2}$ (with fibers $\{\mathfrak{A}_{j}\}_{j\in\overline{\mathbb{Z}_{+}/2}}$) is a $C^{*}$-subalgebra $\mathfrak{A}\subset\prod_{j\in\overline{\mathbb{Z}_{+}/2}}\mathfrak{A}_{j}$ such that: (i) $(\|\mathfrak{a}_{j}\|_{j\in\overline{\mathbb{Z}_{+}/2}})\in C(\overline{\mathbb{Z}_{+}/2})$ for all $(\mathfrak{a}_{j})_{j\in\overline{\mathbb{Z}_{+}/2}}\in\mathfrak{A}$, where $C(\overline{\mathbb{Z}_{+}/2})$ denotes the space of sequences $(\alpha_{j})_{j\in\overline{\mathbb{Z}_{+}/2}}$, $\alpha_{j}\in\mathbb{C}$, such that $\alpha_{\infty}=\lim_{j\to\infty}\alpha_{j}$; (ii) $\alpha\mathfrak{a}\in\mathfrak{A}$ for all $\mathfrak{a}\in\mathfrak{A}$ and $\alpha\in C(\overline{\mathbb{Z}_{+}/2})$. A strict deformation quantization (SDQ) is then defined by the following data:
@@
    We investigate further the relationship between classical and quantum thermal equilibrium with a specific focus on phase transitions. The latter describe the uniqueness/non-uniqueness of KMS states and are of utmost relevance for describing when a physical system undergoes an abrupt change in its macroscopic behaviour, e.g. gas-to-liquid condensation. It is common folklore that classical and quantum phase transition at non-vanishing temperature should be in bijection. Yet, to the best of our knowledge, no mathematically rigorous proof of this claim has been given. In Section 4 we prove that, within the model $\mathbb{S}^{2}$-valued lattice spin system considered in Section 2, under a sufficiently mild assumption, cf. (44), classical and quantum KMS states are unique for temperatures higher than a common sufficiently high threshold temperature. This results is a consequence of Theorems 4.1, 4.7 which provides two new sufficient conditions for the uniqueness of KMS classical and quantum states. Thus, our result is in line with the claimed equivalence between classical and quantum phase transitions.
It is worth to point out that our results are companions of other existing works in this area. In particular, [44] already described an abstract framework which covers the Berezin SDQ for a lattice system in a finite region. Our result (I) generalizes this setting for a specific model but allowing to deal with a spin lattice system on an infinitely extended region: This is important because physically interesting results on thermal equilibrium, e.g. phase transitions, can only be described on infinitely extended system. Concerning (II), in [25] the semiclassical limit of KMS quantum states has been investigated under an assumption which is an abstract version of our Lemma 3.4. In [3] the classical and quantum KMS conditions were related for the case of the Bose-Hubbard system on a finite graph. Similarly, [58] deals with the semiclassical limit of Gibbs quantum and classical states, i.e. KMS states on a spin lattice system associated with a finite region: From this point of view, our result (II) can be seen as a generalization of [58] to a physically more interesting scenario. Finally, concerning (III), Theorems 4.1, 4.7 are inspired by [15, Prop. 6.2.45], which provides a sufficient condition for uniqueness of KMS quantum states with an argument based on a quantum version of the Kirkwood-Salzburg equations. Our result provides a classical analogous of [15, Prop. 6.2.45], moreover, it strengths some of its conclusion —specifically the $j$-dependence of the inverse critical temperature, cf. Remark 4.9. It is also worth to point out that our results cover the regime of high temperatures. For low temperatures it was shown in [11] that, whenever chessboard estimates can be used to prove a phase transition in the classical model, the corresponding quantum model will have a similar phase transition provided $\beta^{2}\ll j$.
@@ Acknowledgements.
Conflict of interest statement.
The authors certify that they have no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter discussed in this manuscript.
@@ 2.2 Berezin SDQ for a single site system
To begin with we consider the Lie group $SU(2)$ and denote by $\{J_{i}\}_{i=1}^{3}$ the generators of the corresponding Lie algebra $\mathfrak{su}(2)$ with commutation relations $[J_{1},J_{2}]=iJ_{3}$ and extended cyclically. Adopting the standard physicist’s notation we denote by $\boxed{D^{(j)}}\colon SU(2)\to M_{2j+1}(\mathbb{C})$ the irreducible representation of $SU(2)$ of spin $j\in\mathbb{Z}_{+}/2$, cf. [29, §5.4]. We will denote by
$|j,m\rangle\in\mathbb{C}^{2j+1}\qquad m\in\{-j,\ldots,j\}\,,$ (10)
@@
$Q_{j}(a):=\int_{\mathbb{S}^{2}}a(\sigma)|j,\sigma\rangle\langle j,\sigma|\,\mathrm{d}\mu_{j}(\sigma)\,.$ (13)
@@
identify a SDQ over a suitably defined bundle of $C^{*}$-algebras $\boxed{B_{\ast}}\subseteq\prod_{j\in\overline{\mathbb{Z}_{+}/2}}B_{j}$, cf. [40, Thm. 8.1]. From a physical point of view, the semiclassical parameter is identified with $\boxed{h_{j}}:=(2j+1)^{-1}$.
@@
where $P_{\ell,m}$ denotes the Legendre polynomial of order $\ell,m$. Notice that this choice of normalization is such that $\|Y_{\ell,m}\|_{L^{2}(\mathbb{S}^{2},\mu_{\ell})}=1$, moreover, $\|Y_{\ell,m}\|_{B_{\infty}}\leq 1$, cf. [55, Cor. 2.9]. The set $\{Y_{\ell,m}\}_{\ell,m}$ is a complete orthogonal system for $L^{2}(\mathbb{S}^{2},\mu_{0})$ made by orthogonal but not $L^{2}$-normalized vectors. With this convention we also have [45, §3.6.2]
@@
where we used Schur orthogonality relations, cf. [37, §4.10], while $\boxed{\textsc{cg}_{j_{1},m_{1};j_{2},m_{2}}^{j,m}}$ denotes the Clebsch-Gordan coefficients, cf. [45, §3]. We recall in particular that $\textsc{cg}_{j_{1},m_{1};j_{2},m_{2}}^{j,m}\in\mathbb{R}$, moreover, the coefficient vanishes unless $m=m_{1}+m_{2}$ and $|j_{1}-j_{2}|\leq j\leq j_{1}+j_{2}$. For later convenience it is also worth recalling that, for all $\sigma\in\mathbb{S}^{2}$,
@@
$\langle A|B\rangle_{\textsc{hs}}:=\frac{1}{2j+1}\operatorname{tr}(A^{*}B)\,,$ (19)
@@
By direct inspection $\tilde{D}^{(j)}(R)$ is unitary with respect to the Hilbert-Schmidt scalar product (19) on $M_{2j+1}(\mathbb{C})$, moreover $R\mapsto\tilde{D}^{(j)}(R)$ is a unitary representation of $SU(2)$. Though $D^{(j)}$ is irreducible, $\tilde{D}^{(j)}$ is not irreducible and by Peter-Weyl Theorem, cf. [29, §5.2], it decomposes into irreducible representations $\{D^{(\ell)}\}_{\ell\in\mathbb{Z}_{+}/2}$. Notably the Berezin quantization map $Q_{j}\colon C(\mathbb{S}^{2})\to M_{2j+1}(\mathbb{C})$ of Equation (13) can be used to provide the explicit decomposition of $\tilde{D}^{(j)}$ in its irreducible components. To this avail, let
@@
where in the last line we used the symmetry property of the Clebsch-Gordan coefficients, cf. [37, §8]. Equations (23)-(25) imply that
@@
form a complete orthogonal system in $M_{2j+1}(\mathbb{C})$ with respect to the Hilbert-Schmidt scalar product (19). (As an aside, we observe that this fact provides a quick proof of that $Q_{j}\colon\dot{B}_{\infty}\to B_{j}$ is surjective: Indeed, for any $A_{j}\in B_{j}$ we may consider $a_{j}:=\sum_{\ell=0}^{2j}\sum_{m=-\ell}^{\ell}A_{\ell,m}Y_{\ell,m}\in\dot{B}_{\infty}$ where $A_{\ell,m}:=\|Q_{j}(Y_{\ell,m})\|_{\textsc{hs}}^{-2}\langle Q_{j}(Y_{\ell,m})|A_{j}\rangle_{\textsc{hs}}$ so that $A_{j}=Q_{j}(a_{j})$.) Moreover, Equation (22) ensures that $Q_{j}$ intertwines between the left-action representation $R\mapsto\hat{R}$ and $\tilde{D}^{(j)}$. Since $R\mapsto\hat{R}$ is unitary equivalent to $D^{(\ell)}$ when restricted to the vector space spanned by $\{Y_{\ell,m}\}_{m\in[-\ell,\ell]\cap\mathbb{Z}}$, it follows that $\tilde{D}^{(j)}$ is unitary equivalent to $D^{(\ell)}$ when restricted to the vector space spanned by $\{Q_{j}(Y_{\ell,m})\}_{m\in[-\ell,\ell]\cap\mathbb{Z}}$. Thus, $\tilde{D}^{(j)}$ is not irreducible and decomposes into direct sum of the irreducible representations of $SU(2)$ with total spin $\ell\in\{0,\ldots,2j\}$, each of which taken with multiplicity one.
@@ 2.3 Berezin SDQ on $\Gamma$
To this avail we recall that in [44] the Berezin SDQ $Q_{j}\colon\dot{B}_{\infty}\to B_{j}$ has been lifted to a SDQ $\boxed{Q_{j}^{\Lambda}}\colon\dot{B}_{\infty}^{\Lambda}\to B_{j}^{\Lambda}$ for any finite region $\Lambda\Subset\Gamma$. In a nutshell, this boils down to define $Q_{j}^{\Lambda}$ by linear extension of the tensor product map $\bigotimes_{x\in\Lambda}Q_{j}$ and checking that the data $B_{\infty}^{\Gamma},B_{\ast}^{\Lambda},\{Q_{j}\}_{j\in\mathbb{Z}_{+}/2}$ fulfil the requirement of a SDQ —here $\boxed{B_{\ast}^{\Lambda}}:=\bigotimes_{x\in\Lambda}B_{\ast}$. A non-trivial task in this setting is to prove that $B_{\ast}^{\Lambda}$ is again a bundle of $C^{*}$-algebras over $\overline{\mathbb{Z}_{+}/2}$: This is addressed in full generality in [44] by using the results of [36].
@@
Let $\{\mathfrak{A}_{j}\}_{j\in\overline{\mathbb{Z}_{+}/2}}$ be a collection of $C^{*}$-algebras. For later convenience we recall the following sufficient condition which identifies a continuous bundle of $C^{*}$-algebras $\mathfrak{A}\subset\prod_{j\in\overline{\mathbb{Z}_{+}/2}}\mathfrak{A}_{j}$ by defining a dense set of (a posteriori) elements of $\mathfrak{A}$cf. [39, Prop. 1.2.3], [40, Prop. C.124]. In more details, let $\dot{\mathfrak{A}}\subseteq\prod_{j\in\overline{\mathbb{Z}_{+}/2}}\mathfrak{A}_{j}$ be such that: